Consider a guitar string that has a tension of 90 N, a length of 65 cm, and a mass per unit length of 7.9 × 10 –4 kg/m. What is the fundamental


1
Expert's answer
2021-02-04T10:19:56-0500

The fundamental frequency is


"f=\frac{1}{2L}\sqrt{\frac T\mu}=260\text{ Hz}."